OpenStudy (anonymous):

On average, a cow and a half can give a pint and a half of milk in 36 hours. How many pints can three cows give on average in 72 hours? (all cows give milk at the same rate.)

OpenStudy (anonymous):

sorry no one seems to be on open study this early so many questions will go unanswered.. still waiting for an answer for mine ;/

OpenStudy (anonymous):

The total work done by 1.5 cows in 36 hours result in 1.5 pints of milk. Since all cows give milk at same rate, we can get the pints of milk (say p) of 3 cows in 72 hours by ratio and proportion...\[\frac{1.5~cows \times 36~hours}{1.5~pints~of~milk}=\frac{3~cows \times72~hours}{p~pints~of~milk}\]\[\frac{\cancel{(1.5)}(36)}{\cancel{1.5}}=\frac{(3)(72)}{p}\]by cross-multiplication, we can have...\[p=\frac{(3)\cancel{(72)}^2}{\cancel{36}}=(3)(2)=6~pints~of~milk\]

OpenStudy (anonymous):

any questions @GizmoGirl04 ?